OpenStudy (anonymous):

Find the exact value of cos 240°.

OpenStudy (austinl):

Use a unit circle, it is really easy.

OpenStudy (anonymous):

well cos 240 = cos(180+60) which is cos(180+x), where x is any acute angle. and cos(180+x) = -(cos x) cos 240 is in quadrant 3 so it will be negative. so cos(180+60) = -(cos 60) = - 1/2,